OpenStudy (anonymous):

3. A regular pentagon has an apothem of 3.2 m and an area of 37.2 m². What is the length of one side of the pentagon? 3.96 m 4.65 m 11.875 m 23.75 m

OpenStudy (anonymous):

Use the formula A = .5(apothem)(perimeter)

OpenStudy (anonymous):

Ok but first, what in the he!! is an apothem

OpenStudy (anonymous):

I dont have a perimeter just an apothem

OpenStudy (anonymous):

37.2m2= .5(3.2)(x)

OpenStudy (anonymous):

@bandicoot12

OpenStudy (anonymous):

sorry about that

OpenStudy (anonymous):

An apothem is kind of like the radius of a shape

OpenStudy (anonymous):

ok well now that its set up what do i do?

OpenStudy (anonymous):

A = 0.5 * Apothem * Perimeter As you said earlier... 37.2m^2= .5(3.2)(x) Solving for x will give you the perimeter not the side length so divide the perimeter by 5 (because it is a regular pentagon) and you will have your answer

OpenStudy (anonymous):

so the peremiter is 1.6

OpenStudy (anonymous):

How did you get 1.6?

OpenStudy (anonymous):

Here is some info on apothems :) http://www.mathopenref.com/apothem.html

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

.5*3.2

OpenStudy (anonymous):

37.2 = .5 * 3.2 * perimeter

OpenStudy (anonymous):

Yeah i've got that down but I don't know where to start that from there

OpenStudy (anonymous):

ok let p represent perimeter 37.2 = .5 * 3.2 * p 37.2 = 1.6 * p (37.2/1.6) = p p = 23.25 What do we do from here?

OpenStudy (anonymous):

Oh see i didn't know what to do from the second step.

OpenStudy (anonymous):

4.565

OpenStudy (anonymous):

yep, good job!

OpenStudy (anonymous):

THank you so much!

OpenStudy (anonymous):

no need to thank me

OpenStudy (anonymous):

hahaha nice ;P